Complete the table below to identify the market value per unit amounts for each product, the total cost amounts, and the total market value amounts. Cost per Market ! O Data Table Group # Smart phone case Nolepad case Total Cost per Unit Total Market Manufacturer General Tech General Tech # of Units Unit Group # Manufacturer CRC#0f Units Sp Disposal NPM Cost General Tech General Tech Pacific Base Pacific Base Subtotal General Tech Smart phone case Notepad case Smart phone case Notepad case 5 00$11 S 650 350 600 1 S 8$5 18 27 10 19 Smart phone case Pacific Base Nolepad case Pacific Base 13 26 18 17 Subtotal Pacific Base Print Done Requirement a. Conduct a lower-ol-cost-or-market test for Case Closed assuming that it uses the individua-Htem approach for LCM computations. (Enter a if no LCM loss adjustment is required Assuming Case Closed uses the individual-item approach for LCM computations, it will record a$ loss adjustment. Requirement b. Conduct a lower-of-cost-or-market test for Case Closed assuming that it uses the group-by-group approach for LCM computations. (Assume the groups are based on the inventory manufacturer. Enter a O if no LCM loss adjustment is required.) Assuming Case Closed uses the group-by-group approach for LCM computations, it will record a $

Requirement c. Prepare the journal entry needed to adjust Case Closed inventory to a​ lower-of-cost-or-market (LCM) basis assuming that LCM is applied to individual items. Case ClosedCase Closed uses the indirect approach to record any adjustments to LCM.
Answer #1
Repalcement cost shall be taken as the market price.But,it is subject to a ceiling and floor limit
Ceiling limit =Estimated selling price-completion and disposal cost
Floor limit=Seiling limit-Normal markup %
a. Compute LCM for each and every item of inventory in individual item approach
Market per unit:
Group # Manufacturer Ceiling limt
For market
price
Floor limt
For market
price

Market
price
(SP-Disposal) [Ceiling limit-NPM] CRC subject
to ceiling and floor limit
Smartphone case General tech 10 2 5
Notepad case General tech 24 14 18
Smartphone case Pacific base 4 1 2
Notepad case Pacific base 23 5 13
LCM computation
Group # Manufacturer # of units Cost per unit Total cost Market
per unit
Total
market
LCM
1 2 Lower of
1 & 2
Smartphone case General tech 400 5 2000 5 2000 2000
Notepad case General tech 650 19 12350 18 11700 11700
Subtotal-General tech 14350 13700
Smartphone case Pacific base 350 5 1750 2 700 700
Notepad case Pacific base 600 17 10200 13 7800 7800
Subtotal-Pacific base 11950 8500
b. Compute LCM for each group of inventory
LCM computation
Group # Manufacturer # of units Cost per unit Total cost Market
per unit
Total
market
LCM
1 2 Lower of
1 & 2
Smartphone case General tech 400 5 2000 5 2000
Notepad case General tech 650 19 12350 18 11700
Subtotal-General tech 14350 13700 13700
Smartphone case Pacific base 350 5 1750 2 700
Notepad case Pacific base 600 17 10200 13 7800
Subtotal-Pacific base 11950 8500 8500
c. Total cost of all groups=14350+11950=$ 26300
Total market value of all groups=13700+8500=$ 22200
Inventory write down required=26300-22200=$ 4100
Account Debit Credit
Loss on inventory write-down 4100
Allowance to reduce inventory to market value 4100